OpenStudy (anonymous):

how much playing area is outside the indicated part? round to the nearest whole number where necessary. info - a professional basketball court, outside the two rectangular free throw lanes each 19ft by 14ft. Professional Basketball Court - L 94ft and W 50ft

OpenStudy (anonymous):

The answer will be the total area of the court minus the area of the indicated parts, where area is given by LxW. So (Playing area outside the indicated parts) = total area - area of indicated parts = (94*50) - 2*(19*14) (because there are two free throw lanes) = 4700 - 532 = 4168 square feet